# Cryptomite: A versatile and user-friendly library of randomness extractors

### v1

@@ A.2 Proofs for the $\mathsf{Circulant}$ extractor
Total Computation Time: The $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor requires two pre-processing steps of $O(n)$: first, to reduce $n$ so that it is a prime with 2 as a primitive root, and second, to check that the input $X$ is not the all 00 or all $1$ string. The extractor itself can be implemented using the NTT, as shown in Algorithm 1. As shown in Theorem 14, this has computation time $O(n\log n)$. Therefore, the total computation time for implementing the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor is $O(n\log n)$ (sometimes called near-linear or quasi-linear).
Circulant extractor implementation – The $\mathsf{Circulant}$ extractor is implemented using the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ implementation above, with two small differences that do not affect the overall computation time. First, to check that the input length plus one (i.e. the length of $x$ plus one) is a prime, instead of prime with primitive root 2 (which is lifted by a change in proof techniques) and second, to pad the $x$ input string with a single (0) bit.
A.4 $\mathsf{Toeplitz}$ implementation
Although the definition of the $\mathsf{Toeplitz}$ extractor is essentially a circular convolution, the $O(n\log n)$ algorithm in Section A.1 expects the input and output lengths of the convolution to be equal, so it cannot be directly applied. Therefore, to implement $\mathsf{Toeplitz}$ in $O(n\log(n))$ time, we must embed the Toeplitz matrix within a square Circulant matrix. Concretely, the Toeplitz matrix can be embedded in the top-left quadrant of the Circulant matrix Equation 101, so we can perform $\mathsf{Toeplitz}$ extraction using the NTT.
$\left(\begin{array}[]{ccccc|ccccc}x_{0}&x_{k-1}&\ldots&\ldots&x_{k-n+1}&0&x_{n-1}&\ldots&x_{2}&x_{1}\\ x_{1}&x_{0}&\ddots&&\vdots&x_{k-n+1}&\ddots&\ddots&&x_{2}\\ x_{2}&\ddots&\ddots&\ddots&\vdots&\vdots&\ddots&&\ddots&\vdots\\ \vdots&&\ddots&x_{0}&x_{k-1}&x_{k-2}&&\ddots&\ddots&x_{n-1}\\ x_{n-1}&\ldots&\ldots&x_{1}&x_{0}&x_{k-1}&x_{k-2}&\ldots&x_{k-n+1}&0\\ \hline\cr 0&x_{n-1}&\ldots&\ldots&x_{1}&x_{0}&x_{k-1}&\ldots&\ldots&x_{k-n+1}\\ x_{k-n+1}&\ddots&\ddots&&x_{2}&x_{1}&x_{0}&\ddots&&\vdots\\ \vdots&\ddots&&\ddots&\vdots&x_{2}&\ddots&\ddots&\ddots&\vdots\\ x_{k-2}&&\ddots&\ddots&x_{n-1}&\vdots&&\ddots&x_{0}&x_{k-1}\\ x_{k-1}&x_{k-2}&\ldots&\ldots&0&x_{n-1}&\ldots&\ldots&x_{1}&x_{0}\end{array}\right)$ (101)
The full implementation of the $\mathsf{Toeplitz}$ extractor can be seen in the pseudo-code in Algorithm 2.
@@ A.4 $\mathsf{Toeplitz}$ implementation
[…]
2:     $L\leftarrow 2^{\text{floor}(\log_{2}(2n)+1)}$triangleright Calculate NTT size
@@
A.5 $\mathsf{Trevisan}$ implementation
We follow the implementations by Mauerer et al. [19], which combines multiple iterations of the weak design from Hartman and Raz [50] to create a block weak design, then uses the strong $\mathsf{RSH}$ polynomial hashing based one-bit extractor constructed from a combination of the Reed-Solomon and the Hadamard code (see [19], ‘Section III. DERIVATIONS’). Informally, our implementation of the $\mathsf{Trevisan}$ extractor takes an input $x\in\{0,1\}^{n}$ and a seed $y\in\{0,1\}^{d}$ and extracts $m$ output bits by
@@ A.5 $\mathsf{Trevisan}$ implementation
    Using each of these $m$ sub-string to (respectively) seed a $\mathsf{RSH}$ one-bit one-bit extractor, processes the extractor input $x$ into a near-perfect bit as output.
@@
$\mathsf{RSH}$ 1-bit extraction: After using the block weak design to split the seed into $m$ sub-strings, we perform $m$ Reed-Solomon Hadamard (called $\mathsf{RSH}$) one-bit extractions, where the one bit extractor is constructed from the concatenation of a Reed-Solomon and a Hadamard code (see [19], ‘Section III. C. 3. Polynomial hashing’). Each one-bit extractors takes as input the extractor input $x\in\{0,1\}^{n}$ and one of the $m$ sub-strings of the seed $y$ generated by the block weak design, $y_{S_{i}}\in\{0,1\}^{t}$ for some $i\in\{1,\ldots,m\}$. To implement $\mathsf{RSH}$, the seed sub-string $y_{S_{i}}$ is further split into two sub-strings of $l$ bits, where the first is used in the Reed-Solomon step and the second in the Hadamard step. The initial $\mathsf{RSH}$ seed is of length $t\geq 2l$, where $l$ is
@@
  • Split the extractor input $x$ into $s=\lceil n/l\rceil$ chunks $x_{i}$. Pad the last chunk $x_{s}$ with 00s if need be to make all chunks of length $l$.
@@
A.6 $\mathsf{Von\hskip 2.84544ptNeumann}$ implementation

### v3

@@ A.3 Proofs for the $\mathsf{Circulant}$ extractor
Total Computation Time: The $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor requires two pre-processing steps of $O(n)$: first, to reduce $n$ so that it is a prime with 2 as a primitive root, and second, to check that the input $X$ is not the all 0 or all $1$ string. The extractor itself can be implemented using the NTT, as shown in Algorithm 1. As shown in Theorem 14, this has computation time $O(n\log n)$. Therefore, the total computation time for implementing the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor is $O(n\log n)$ (sometimes called near-linear or quasi-linear).
Circulant extractor implementation – The $\mathsf{Circulant}$ extractor uses the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ implementation described above, with one minor difference that does not affect the overall computation time: the input $x$ is padded with a single 0 bit.
A.5 $\mathsf{Toeplitz}$ implementation
Although the definition of the $\mathsf{Toeplitz}$ extractor is essentially a circular convolution, the $O(n\log n)$ algorithm in Section A.2 expects the input and output lengths of the convolution to be equal, so it cannot be directly applied. Therefore, to implement $\mathsf{Toeplitz}$ in $O(n\log(n))$ time, we must embed the Toeplitz matrix from Equation 33, with $m$ rows and $n$ columns, within a square circulant matrix. Concretely, the Toeplitz matrix is embedded in the top-left quadrant of the square $(n+m-1)\times(n+m-1)$ circulant matrix Equation 63:
$\mathrm{circ}(y_{0},y_{n+m-2},y_{n+m-3},\ldots,y_{1})=\begin{bmatrix}\mathrm{toep}(y)&\cdot\\ \cdot&\cdot\end{bmatrix}$, (63)
where $\mathrm{circ}(\cdot)$ denotes the circulant matrix generating function as in Equation 24 and $\mathrm{toep}(y)$ denotes the Toeplitz matrix from Equation 33. Then, by padding the weak input $x\in\{0,1\}^{n}$ with m1m-1 zeros (i.e., $x\to[x,\{0\}^{m-1}]$), the $\mathsf{Toeplitz}$ extractor can be recovered by the matrix-vector multiplication $(\mathrm{circ}(y_{0},y_{n+m-2},y_{n+m-3},\ldots,y_{1})x^{\prime})_{0:m-1}$ where $x^{\prime}=[x,\{0\}^{m-1}]$. Using this embedding, the $\mathsf{Toeplitz}$ extractor can be implemented using the NTT in $O(n\log n)$ computation time. The full implementation of the $\mathsf{Toeplitz}$ extractor can be seen in the pseudo-code in Algorithm 2.
@@ A.5 $\mathsf{Toeplitz}$ implementation
1:function ExtToeplitzn,m(u, v)
2:  $L\leftarrow 2^{\text{floor}(\log_{2}(2n)+1)}$ $\triangleright$ Calculate NTT size
@@
A.6 $\mathsf{Trevisan}$ implementation
We follow the implementations by Mauerer et al. [21], which combines multiple iterations of the weak design from Hartman and Raz [61] to create a block weak design, then uses the strong Reed-Solomon Hadamard (denoted $\mathsf{RSH}$) polynomial hashing based one-bit extractor constructed from a combination of the Reed-Solomon and the Hadamard code (see [21], ‘Section III. DERIVATIONS’). Informally, our implementation of the $\mathsf{Trevisan}$ extractor takes an input $x\in\{0,1\}^{n}$ and a seed $y\in\{0,1\}^{d}$ and extracts $m$ output bits by
@@ A.6 $\mathsf{Trevisan}$ implementation
    Using each of these $m$ sub-string to (respectively) seed a $\mathsf{RSH}$ one-bit one-bit extractor, process the extractor input $x$ into a near-perfect bit as output.
@@
$\mathsf{RSH}$ 1-bit extraction: After using the block weak design to split the seed into $m$ sub-strings, we perform $m$ $\mathsf{RSH}$ one-bit extractions, where the one bit extractor is constructed from the concatenation of a Reed-Solomon and a Hadamard code (see [21], ‘Section III. C. 3. Polynomial hashing’). Each one-bit extractor takes as input the extractor input $x\in\{0,1\}^{n}$ and one of the $m$ sub-strings of the seed $y$ generated by the block weak design, $y_{S_{i}}\in\{0,1\}^{t}$ for some $i\in\{1,\ldots,m\}$. To implement $\mathsf{RSH}$, the seed sub-string $y_{S_{i}}$ is further split into two sub-strings of $l$ bits, where the first is used in the Reed-Solomon step and the second in the Hadamard step. The initial $\mathsf{RSH}$ seed is of length $t\geq 2l$, where $l$ is
@@
  • Split the extractor input $x$ into $s=\lceil n/l\rceil$ chunks $x_{i}$. Pad the last chunk $x_{s}$ with 0s if need be to make all chunks of length $l$.
@@
A.7 $\mathsf{Von\hskip 2.84544ptNeumann}$ implementation